OpenStudy (anonymous):

-x + 2y + 8x - 5y + (-6x) - -4y

OpenStudy (anonymous):

-x + 8x - 6x = x 2y-5y+4y= y

OpenStudy (anonymous):

thanks:)

OpenStudy (anonymous):

-x +2y +8x -5y +(-6x) --4y First, +(-6x) = -6x, and --4y = +4y. So that yields: -x +2y +8x -5y -6x +4y Now group the Xs and the Ys: +8x -6x -x = x +2y +4y -5y = y so -x +2y +8x -5y +(-6x) --4y = x+y